Circular No. 7367

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

COMETS C/1999 O1 AND C/1999 P3 D. A. Biesecker, SM&A Corporation and Goddard Space Flight Center,

reports measurements of two apparent Kreutz sungrazing tailless comets (discovery observations below) both tailless) observed with SOHO’s LASCO C3 coronagraph; astrometry and orbital elements appear on MPEC 2000-C59 and 2000-C60. C/1999 O1 was found by T. Lovejoy in movies posted at the SOHO website. C/1999 P3 was found by D. Lewis.

1999 UT α_{2000} δ_{2000} Comet
July 31.513 8^{h}27.^{m}0 +16^{o}48′  C/1999 O1
Aug. 3.238 8 33.8 +15 58  C/1999 P3

NOVA SAGITTARII 2000 N. M. Ashok, U. S. Kamath, T. Chandrasekhar, and Anandmayee Tej,

Physical Research Laboratory, Ahmedabad, report 1.1–2.5-μm spectrophotometry of N Sgr 2000 using the 1.2-m telescope (+ PRL Near Infrared Imaging Spectrograph + HgCdTe NICMOS3 detector array) at Mt. Abu Infrared Observatory on Feb. 12.00 UT. The JHK-band spectra show strong H I emission lines (FWHM ∼ 1900 km/s) at Brγ, Br_{10} to Br_{15}, Paβ, and Paγ. The Lyβ-fluoresced O I line at 1.1287 μm is seen with significant strength. Also seen are the N I line at 1.2461 μm and O I at 1.3164 μm.

SUPERNOVA 2000H IN IC 454 A. Pastorello, G. Altavilla, E. Cappellaro, and M. Turatto, Osservato-

rio Astronomico di Padova; and S. Benetti, Telescopio Nazionale Galileo, obtained a spectrum (range 350–750 nm, resolution 2 nm) of SN 2000H on Feb. 10.98 UT with the Asiago 1.82-m telescope (+ AFOSC), with different conclusions from Garnavich et al. (IAUC 7366): “The low-S/N spectrum shows that the supernova is very similar to the type-Ic SN 1996aq (cf. IAUC 6454) around maximum. The strongest features are P-Cyg lines of Ca II (absorption measured at 385 nm), Fe II (447, 489, 507 nm), and Na I D; the redshift of the parent galaxy reported in NED is 3945 km/s. Also present is an enigmatic broad, asymmetric absorption with a minimum at 637 nm. Guided by the similarity with the type-Ic SN 1987M (Filippenko 1992, Ap.J. 384, 37L), it is suggestive to attribute this feature to a blend of Si and H, though in the case of SN 2000H, the feature is unusually strong. Following the evolution of SN 2000H through the nebular phase could give useful insights on the connection between supernova types Ic and II.”

2000 February 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green